OpenStudy (anonymous):

y is > or less than 0.5x^2-5. show me how to do it please in vertex form

OpenStudy (mertsj):

Do it like Satellite showed you. Or if you know calculus, do it like Loser shower you.

OpenStudy (anonymous):

i never got satellite's relpy:/ unless i just need to refresh the page

OpenStudy (mertsj):

first coordinate of the vertex is −b/2a which in your case is −(−2)/2=1

OpenStudy (mertsj):

satellite73 Medals 0 second coordinate of the vertex is what you get when you replace x by 1

OpenStudy (mertsj):

satellite73 Medals 0 to write it in "vertex form" put x^2−2x−10=(x−1)^2+k where k is the second coordinate of the vertex found above

OpenStudy (mertsj):

That is what Satellite showed you to do.